OpenStudy (anonymous):

can you give me the best tips on how i can get the LCD of a rational expressions in adding and subtracting rational expressions?

OpenStudy (anonymous):

what is LCD ?

OpenStudy (anonymous):

Least common denominator

OpenStudy (anonymous):

can you give an example of what you're trying to do ? you're question seem a bit odd to me.

OpenStudy (anonymous):

for example\[6a/a^2-b^2 +2a/ab+b^2 - 3/a-b \]

OpenStudy (anonymous):

the denominators are \[a^2-b^2, ab+b^2 , a-b\]

OpenStudy (anonymous):

herm, you should add parenthesis then, because i can't see what's what

OpenStudy (anonymous):

for example i think it's more 6a/(a^2 - b^2) + 2a/(ab + b^2) -3/(a-b)

OpenStudy (anonymous):

you want to factor the denominators. See what they have in common and what they are missing: \[a^2-b^2 = (a+b)(a-b)\] \[ab+b^2 = b(a+b)\] and a - b stays the same. Then multiply the numerator and denominator by whats missing. looking at the first denominator, we need a 'b' term there, so multiply top and bottom by b. The second denominator, we are missing a 'b' and a '(a-b)' Third denominator is missing a 'b' and a '(a+b)'

OpenStudy (anonymous):

Say u have a number of terms all with different denominators. Start at the first, include that. Go to the second, include that it if it is different to the first. Go to the third, include it if it is different to the first two etc eg 1/a + 1/b + 1/c LCD is abc.

OpenStudy (anonymous):

I just made that up, see if u can find an example where it doesn't work.

OpenStudy (anonymous):

It's a bit more complicated if we are talking about actual numbers rather than algebra.